In a department store there are 120 customers, 90 of who will buy at least 1 item. If 5 customers are selected at random, one by one, find the probability that all will buy at least 1 item.
Answer by stanbon(75887)

In a department store there are 120 customers, 90 of who will buy at least 1 item. If 5 customers are selected at random, one by one, find the probability that all will buy at least 1 item.
-------
P(each of 5 buy at least one) = 1 - P(none of 5 buy at least one)
---
Note: There are 30 who will not buy at least one item
---------------
= 1 - 30C5/120C5
----
= 1 - 0.0000007478
---
= 0.999252....
======================
Cheers,
Stan H.